.66x-.75y=-8
.33x+.5y=17
How do you solve by substitution?
Show work?

Answer by checkley75(3666)

.66x-.75y=-8
.33x+.5y=17
First you need to represent eith the x or the y unknown in terms of the other unknown.
.33x=17-.5y
x=(17-.5y)/.33
Now replace x in the other equation with (17-5y)/.33 & solve for y.
.66(17-.5y)/.33-.75y=-8
2(17-.5y)-.75y=-8
34-y-.75y=-8
-1.75y=-8-34
-1.75y=-42
y=-42/-1.75
y=24 ans.
x=(17-.5*24)/.33
x=(17-12)/.33
x=5/.33
x=15.1515